OpenStudy (anonymous):

use the properties of logarithms to expand as a sum, difference, or constant. (all variables are positive) log 5X 10

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

\[\log(AB)=\log(A)+\log(B)\] is what you need

OpenStudy (anonymous):

so \[\log(5x)=\log(5)+\log(x)\]

OpenStudy (anonymous):

what does it mean by "sum, difference, or constant"

OpenStudy (anonymous):

"sum" as in addition

OpenStudy (anonymous):

i.e. \(\log(5)+\log(x)\) is a sum

OpenStudy (anonymous):

oh ok, i get it now. I took notes on those today. I just have no idea on how to use them.